OpenStudy (anonymous):

The Caribbean paradise has 28 more deluxe staterooms than the pacific paradise. The Mediterranean paradise has 40 fewer deluxe staterooms than three times the number of deluxe staterooms on the pacific paradise. Find the number of deluxe staterooms for each ships if the total number of deluxe staterooms for the three ships is 793?

OpenStudy (anonymous):

this is how far I got. P+C+M=793 AFTER THAT I AM STUCK

OpenStudy (texaschic101):

P + C + M = 793 C = P + 28 M = 3P - 40 Now just sub in C and M in the first equation and that will get you P. Then sub that into the other 2 equations and that will get you C and M.

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

if you need anymore help with it, just tag me :)

OpenStudy (anonymous):

P+C+M=793 C=P+28 M=3P-240 P+P+P+28-80

OpenStudy (texaschic101):

P + C + M = 793 P + (P + 28) + (3P - 40) = 793 -- combine like terms 5P - 12 = 793 -- add 12 5P = 793 + 12 5P = 805 -- divide by 5 P = 805/5 P = 161 Does that help ?

OpenStudy (anonymous):

Yes thank you

OpenStudy (texaschic101):

no problem :)